OpenStudy (anonymous):

i...just....cant....do it. lim x->4 of (4x-x^2)/(2-sqrt(x))

OpenStudy (campbell_st):

I think you'll find the limit is undefined... no matter what you do the result will be a zero in the denominator.

OpenStudy (anonymous):

erm. when i put it in a limit calculator i get 16

OpenStudy (anonymous):

im @ \[\lim_{x \rightarrow 4} (4x-x^2)/(2-\sqrt{x)} * ((2+\sqrt{x})/(2+\sqrt{x}))\]

OpenStudy (anonymous):

= (8x-2x^2+4xrt(x)-x^2rt(x)) / 2-x

OpenStudy (anonymous):